OpenStudy (phillyboy_drew):

What are the zeros of the function h(x) = x³ – x + 6?

OpenStudy (welshfella):

thers no standard way to solve cubic equations but as 2^3 = 8 what do you think one zero will be?

OpenStudy (welshfella):

oh sorry i should have said (-2)^3 = -8 try replacing x by -2 in the equation and see what you get

OpenStudy (welshfella):

(-2)^2 - (-2) + 6 - what does that work out to?

OpenStudy (welshfella):

if that works out to 0 if you have found one zero (-2)

OpenStudy (radar):

Factor the cubic getting (x+2)(x^2 - 2x + 3) set to 0 now each factor can be solved, note that one zero is as welshfella said, -2 you need to solve the quadratic factor x^2 -2x + 3 = 0. This will solve the other 2 complex zeros